OpenStudy (anonymous):

What would you need to do to find the intersection of y=e^(2x) and x=ln(3)?

OpenStudy (anonymous):

y=e^(2x) y=e^(2ln3) y=e^(ln3^2) y=e^(ln9) y=9

OpenStudy (anonymous):

Oy, very clever! Thanks.

OpenStudy (anonymous):

So, point of intersection is (ln3 , 9)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yeah, thank you, Saurav.

OpenStudy (anonymous):

welcome